OpenStudy (anonymous):

Write the quadratic function in vertex form. y = x2 - 14x + 53 A. y = (x - 7)2 - 4 B. y = (x + 7)2 - 4 C. y = (x - 7)2 + 4 D. y = (x + 7)2 + 4

OpenStudy (anonymous):

i cant figure this out..

OpenStudy (campbell_st):

rewrite the equation y - 53 = x^2 -14x + ? you now need to complete the square

OpenStudy (anonymous):

what do you mean complete the square? i'm horrible at math..im sorry

OpenStudy (campbell_st):

this is a perfect square \[(x - a)^2 = x^2 - 2ax + a^2\] you need to halve the middle term... -14 then square it... add the value to both sides of the equation... thats the process of completing the square... you are looking to find a so you can calculate a^2

OpenStudy (anonymous):

so would the answer be.. B. y = (x + 7)2 - 4

OpenStudy (campbell_st):

not quite... what is half of -14..?

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

ohhh..so id be C

OpenStudy (campbell_st):

ok... so you will have to add (-7)^2 to both sides y - 53 + 49 = x^2 - 14x + 49 the write hand side can be written as a perfect square \[y - 4 = (x - 7)^2\] all you need to do now is add 4 to both sides

OpenStudy (campbell_st):

lol... right hand side... not write hand side... sorry about that

OpenStudy (anonymous):

but then what answer choice would that be.. ? bc its cancel out the 4 of the left and its be y=(x-7)^2+4 so would it be C. y = (x - 7)2 + 4????

OpenStudy (campbell_st):

ok... in the notation for you choices... )2 and )^2 both mean squared...

OpenStudy (campbell_st):

and it is C... well done

OpenStudy (anonymous):

ok:) thankyou!!